Please help me with this, I would be very grateful :)
9a^2+25b^2

Answer by rothauserc(4718)

9a^2 +25b^2 is an elliptic paraoloid that opens upward
integer solution is a=0, b=0
*************************************************************
roots for variable b
Solve for b:
9a^2+25b^2 = 0
:
Subtract 9 a^2 from both sides:
25 b^2 = -9 a^2
:
Divide both sides by 25:
b^2 = -(9 a^2)/25
:
Take the square root of both sides:
b = 1/5(3i)a
or b = 1/5(-3i)a
Note that when we take the square root of - we get i
*************************************************************
roots for variable a
9a^2 = -25b^2
a^2 = -25b^2/9
a = 1/3(5i)b
or a = 1/3(-5i)b
*************************************************************